OpenStudy (anonymous):

Solve for x: |3x + 12| = 18 x = 2, x = −10 x = 2, x = −2 x = 10, x = −10 x = −2, x = 10

OpenStudy (cwrw238):

the expression in the modulus signs can be negative or positive so you have to solve 2 equations 3x + 12 = 18 and 3x + 12 = -18

OpenStudy (anonymous):

i know that x can equal 2

OpenStudy (cwrw238):

right - thas the solution of first equation

OpenStudy (cwrw238):

now solve the 2nd

OpenStudy (anonymous):

thats where i get stuck

OpenStudy (cwrw238):

3x + 12 = -18 subtract -12 from both sides whatt does that give you/

OpenStudy (anonymous):

oh now i understand since it has an absolute value when it's an odd it becomes positive so that means its 2 and -10

OpenStudy (cwrw238):

2 and -10 is correct

OpenStudy (anonymous):

how about this one Solve for x: 2|2x − 2| + 4 = 20 x = 5, x = −3 x = 5, x = −5 x = 3, x = −3 x = −5, x = 3